Determine the electric field intensity at a point P midway between two point charges of +2.0x10^-7 C and -2.0x10^-7 C if they are 100 cm apart.


1
Expert's answer
2021-05-13T09:21:34-0400

Answer

Both the charges are same in magnitude and apposite in nature. So electric field at midway between charges will be in same direction and equal in magnitude. So

"E_{net}=\frac{Kq}{r^2}+\frac{Kq}{r^2}"


"E_{net}=\frac{2Kq}{r^2}"


"E_{net}=\frac{2(9\times10^9)(2\times10^{-7})}{(0.5)^2}=144\times10^{2}N"